OpenStudy (anonymous):

Please help! In a large bag of Skittles®, each of the 5 colors (red, orange, yellow, blue, green) occurs with the same probability. You reach in and select 2 candies. Find the probability that exactly one of the candies is orange. I'm just checking, the answer is 2/5 or 0.4? Right???

OpenStudy (anonymous):

exactly one out of two means one orange and not orange. orange has probability \(\frac{1}{5}\) not orange has probability \(\frac{4}{5}\) you need to multiply these probabilities together

OpenStudy (ganpat):

so it can be, for orange candy probability will be 1/5, while for any other candy it will be 4/5.. so the probability of event will be = 1/5 * 4/5 = 4/25 = 0.16... i am not sure on this.. let some one confirm this..

OpenStudy (anonymous):

@Ganpat is correct, it is \(\frac{1}{5}\times \frac{4}{5}=\frac{4}{25}\)

OpenStudy (anonymous):

okay, thank so much! :)

OpenStudy (anonymous):

hmmm they said the answer is 0.32